Find the area of a triangle defined by R(0,0), S(6,8), and T(10,0).
Answer by nyc_function(2741)

After plotting each point on the coordinate plane, I found the height to be 8 and the base to be 10.

Area of triangle = 1/2 * base * height.


A = 1/2 * 10 * 8

A = 1/2 * 80

A = 40